What quantity of 75 per cent acid solution must be mixed with a 30 percent solution to produce 1215 mL of a 50 per cent solution?
Answer by checkley77(12844)

.75x+.30(1215-x)=.50*1215
.75x+364.5-.30x=607.5
.75x-.30x=607.5-364.5
.45x=243
x=243/.45
x=540 ml of 75% acid.
1215-540=675 ml of 30% acid.
Proof:
.75*540+.30*675=.50*1215
405+202.5=607.5
607.5=607.5